# Performance of Practical Quantum Oblivious Key Distribution

### v2

@@ IV.2 Security against dishonest sender
To guarantee that Alice will not be able to obtain information about the value of $c$ during the string separation phase, it is necessary to show that Alice does not have access to the information of Bob’s bases choices $\theta^{B}_{I_{0},I_{1}}$ from the commitments sent by Bob during Step (4) of the protocol. As shown by Lemma IV.2, the shared state of the parties after the commitment information is sent is computationally indistinguishable from a state where Alice’s information is independent of $\theta^{B}_{\bar{I}_{t}}$.
Lemma IV.2.
@@
At Step (8) of the protocol, Alice sends Bob the system $\Theta^{A}_{\bar{I}_{t}}$ intended to have the information of her measurement bases. Bob then is able to determine the indices for which $\theta^{A}_{\bar{I}_{t}}$ and $\theta^{B}_{\bar{I}_{t}}$ coincide. With this information, he randomly selects sets $I_{0},I_{1}\in\bar{I}_{t}$ of size $N_{\textnormal{raw}}$ for which all indices are associated with matching (for $I_{0}$) or nonmatching (for $I_{1}$) bases. Then he computes $(J_{0},J_{1})=(I_{c},I_{\bar{c}})$, by flipping the order if the pair $(I_{0},I_{1})$ depending on the value of $c$. Clearly, $(J_{0},J_{1})$ depend on both $\theta^{B}_{\bar{I}_{t}}$ and $c$, but as Lemma IV.3 states, any correlation between $(J_{0},J_{1})$, $c$, and Alice’s information disappears if one does not have access to $\theta^{B}_{\bar{I}_{t}}$.
Lemma IV.3.
@@
A proof of both Lemmas IV.2 and IV.3 can be found in Appendix A.3. By setting $J=I_{t}$, Lemma IV.2 guarantees that Alice’s system’s state after the opening information has been sent is computationally indistinguishable from one that is completely uncorrelated with Bob’s measurement basis information in $\bar{I}_{t}$.
Additionally, by recalling that the value of $c$ is sampled independently of any of the considered systems, we know that the state $\rho_{{A^{\prime},\Theta^{A}_{\bar{I}_{t}},\Theta^{B}_{\bar{I}_{t}}},C}$ before $(J_{0},J_{1})$ is computed has the required product form and, from Lemma IV.3, we conclude that the state of all of Alice’s system at this point is computationally indistinguishable from a state uncorrelated with $C$. Let $\mathcal{E}$ be the operation Alice performs in her system from here to the end the protocol. By using Lemma II.2 (4) and grouping all of Alice’s systems into $S$, we obtain the desired result:
@@
IV.3 Security against dishonest receiver
@@
We consider now the scenario in which Alice runs the protocol honestly and Bob runs an arbitrary program. We want to show that the state after finishing the protocol successfully satisfies Eq. (19). This means that the state at the end of the protocol can be described as a mixture of states where Bob’s system is uncorrelated with at least one of the two strings outputted by Alice. Similarly to the dishonest sender’s case, we assume that Bob’s laboratory consists of everything outside Alice’s, which means that he controls the communication channels and the environment. However, we do not assume that Bob is restricted to efficient computations.
@@
The values of Alice’s output strings depend on several quantities: Alice’s measurement outcomes, the choice of the $I_{t},J_{0},J_{1}$ subsets, and the choice of hashing function $f$ during the post-processing phase of the protocol. From all of these, the only ones that are not made explicitly public during the protocol’s execution are Alice’s measurement outcomes. Instead, partial information of these outcomes is revealed at different steps of the protocol. Let $x^{A}_{J_{0}},x^{A}_{J_{1}}$ be the sub-strings of measurement outcomes used to compute Alice’s outputs $m_{0},m_{1}$, respectively, and let $R$ denote Bob’s system at the end of the protocol (which includes all the systems that Alice sent during the execution of the protocol). In order to prove security we need to show that the joint state of the system $X^{A}_{J_{0}},X^{A}_{J_{1}},R$ can be written as a mixture of states $\rho^{b}$ (with $b\in\{0,1\}$) such that the conditional min-entropy $H_{\textnormal{min}}^{\varepsilon}(X^{A}_{J_{b}}|R)_{\rho^{b}}$ is high enough, so that we can use the leftover hash Lemma II.4 to guarantee that the outcome of the universal hashing $m_{b}=f(x^{A}_{J_{b}})$ is uncorrelated with $R$.
@@
At the start of the protocol the parties share a completely correlated entangled system. If the parties make measurements as intended, their outcomes will be only partially correlated, but if Bob was able to postpone his measurement until after Alice’s reveals her measurement bases, Bob could potentially obtain the whole information of $x^{A}$ by measuring in the appropriate basis on his system. To prevent this, Bob is required to commit his measurement bases and results to Alice before knowing which set is going to be tested. Then a statistical test is performed in Step (7) to estimate the correlation of Alice’s measurement outcomes with with the ones that Bob committed. As Lemma IV.4 states, any state passing the aforementioned test is such that, regardless of how Bob defines the sets $(J_{0},J_{1})$ during the string separation phase, there is a minimum of uncertainty that he has with respect to Alice’s measurement outcomes. Recall that, when Alice is honest, the overall state of the protocol before Step (8) will be a partially classical state, which could be written as a mixture over all of Alice’s classical information. Let $\vec{\tau}=(x^{A}_{I_{t}},\theta^{A},r,\textnormal{com},I_{t},I_{s},\textnormal{open}_{I_{t}})$ denote the transcript of the protocol up to Step (8), and let $\rho_{X^{A}B}(\vec{\tau},J_{0},J_{1})$ be the joint state of Alice’s measurement outcomes and Bob’s laboratory conditioned to $\vec{\tau},J_{0},J_{1}$.
Lemma IV.4.
@@
To reach the desired result, we will first show that a state $\tilde{\rho}_{T,\textnormal{SEP},X^{A},B}$ satisfying Lemma IV.4 (1) also satisfies a tighter version Lemma III.2, and then use Lemma IV.4 (2) to attain the bound for the real protocol’s outcome. Since $\tilde{\rho}_{T,\textnormal{SEP},X^{A},B}$ is classical in both $T$ and SEP we can write the state of the joint system of Alice’s measurement outcomes and Bob’s laboratory as a mixture over all the possible transcripts at that point, that is:
@@
where $P(\vec{\tau},J_{0},J_{1})$ defines a probability distribution which is dependent on Bob’s behavior during the previous steps. We can now separate the $\tilde{\rho}_{X^{A}B}(\vec{\tau},J_{0},J_{1})$ in two categories depending on which of the $x^{A}_{J_{0}},x^{A}_{J_{1}}$ is the least correlated with Bob’s system. Consider the function $b(\vec{\tau},J_{0},J_{1})$ to be equal to 00 if $H_{\textnormal{min}}^{\varepsilon}(X^{A}_{J_{0}}|X^{A}_{J_{1}}B)_{\rho(\vec{\tau},J_{0},J_{1})}\geq H_{\textnormal{min}}^{\varepsilon}(X^{A}_{J_{1}}|X^{A}_{J_{0}}B)_{\rho(\vec{\tau},J_{0},J_{1})}$, and equal to $1$ otherwise. By regrouping the terms from (45) for which the value of $b$ is the same, we can rewrite the joint state as:
@@
where, from Lemma IV.4 and recalling that, as Lemma II.3 (5) states, the min-entropy of a mixture is lower bounded by that of the term with the least min-entropy, we know that
@@
where the second inequality follows from Lemma II.3 (3) and (4), and the max entropy term $H_{\max}(S_{b})$ is upper bounded by the size in bits of the syndrome $\ell=N_{\text{raw}}(f\cdot h(p_{\text{max}}+\delta_{1}))$. Now we can apply Equation (IV.3) to Lemma II.4, which states that, for the outcomes $M_{0},M_{1}$ of the universal hashing by Alice in Step (12) and the Bob’s final system $R$ it holds that
@@
Finally, by applying Lemma II.1 (3) and (4) to Equations (46) and  (49), and then Lemma II.1 (1) to Eq. (44) we get the desired result:
@@
V Experimental Implementation
V.1 Description of the Setup
@@
A schematic representation of the experimental setup can be seen in Fig. 5. Spontaneous parametric down conversion (SPDC), attributed to Alice, is used to create polarization entangled photon pairs in the state ${|{\Psi^{+}}\rangle}=\frac{1}{\sqrt{2}}\left({|{HH}\rangle}+{|{VV}\rangle}\right)$, which are coupled into optical fiber. One photon is sent through a 50/50 fiber beam splitter, probabilistically routing it to one of two polarization projection stages. There, a quarter-wave plate (QWP), a half-wave plate (HWP) and a polarizing beam splitter (PBS) are used to project the photons state onto the linear ($H/V$) or diagonal (+/+/-+ / -) basis, respectively. All photons are sent to superconducting nanowire single photon detectors (SNSPDs) and their arrival time is recorded using a time tagging module (TTM). The second photon of the state ${|{\Psi^{+}}\rangle}$, attributed to Bob, travels through an equivalent probabilistic projection setup.
@@
V.2 Practical protocol
The protocol is identical to that from $\pi_{\textnormal{ROT}}$ as described in Section III, with the following amendments:
@@ V.2 Practical protocol
      $1$ click: assign the correct measured bit value and report a successful round
@@
      $2$ clicks from the same basis: assign a random bit value to the measurement result and report a successful round
@@ V.3 Practical security
Note that the statistical check performed by Alice in the second step of the amended protocol (Section V.2) ensures security under the assumption that there is no coherence in the photon-number basis. This is the case in our implementation, since SPDC produces states of the form $\sum_{n=0}^{\infty}\sqrt{c_{n}}{|{n}\rangle}_{1}{|{n}\rangle}_{2}$ in the number basis $\{{|{n}\rangle}\}$ [48], leaving the individual subsystems in incoherent mixtures of the form $\sum_{n=0}^{\infty}c_{n}{|{n}\rangle}{\langle{n}|}$.
To account for the leakage caused by undetected multi-photon emissions to our OT rate expression, we effectively grant Bob an amount of information about Alice’s measurement outcomes equal to the number of indices in $I_{\bar{I}_{t}}$ associated to multi-photon events, upper bounded by $p_{\textnormal{multi}}(1-\alpha)N_{0}$ for large $N_{0}$. Subtracting this leak to the total entropy expression in Eq. (IV.3) leads to a version of Lemma III.2 for security against dishonest receiver corrected for the experimental implementation, which differs from the theoretical version by replacing Eq (29) with
@@
$\quad\quad+\frac{1}{2}\cdot 2^{n-N_{\textnormal{raw}}\left(\frac{1}{2}-\frac{2\delta_{2}}{1-2\delta_{2}}-h\left(\frac{p_{\textnormal{max}}+\delta_{1}}{\frac{1}{2}-\delta_{2}}\right)-f\cdot h(p_{\textnormal{max}}+\delta_{1})-\frac{p_{\textnormal{multi}}}{\frac{1}{2}-\delta_{2}}\right)},$
VI Discussion

### v3

@@ 4.2 Security against dishonest sender
To guarantee that Alice will not be able to obtain information about the value of $c$ during the string separation phase, it is necessary to show that Alice does not have access to the information of Bob’s bases choices $\theta^{B}_{I_{0},I_{1}}$ from the commitments sent by Bob during Step (4) of the protocol. As shown by Lemma 4.2, the shared state of the parties after the commitment information is sent is computationally indistinguishable from a state where Alice’s information is independent of $\theta^{B}_{\bar{I}_{t}}$.
Lemma 4.2.
@@
At Step (8) of the protocol, Alice sends Bob the system $\Theta^{A}_{\bar{I}_{t}}$ intended to have the information of her measurement bases. Bob then is able to determine the indices for which $\theta^{A}_{\bar{I}_{t}}$ and $\theta^{B}_{\bar{I}_{t}}$ coincide. With this information, he randomly selects sets $I_{0},I_{1}\in\bar{I}_{t}$ of size $N_{\textnormal{raw}}$ for which all indices are associated with matching (for $I_{0}$) or nonmatching (for $I_{1}$) bases. Then he computes $(J_{0},J_{1})=(I_{c},I_{\bar{c}})$, by flipping the order if the pair $(I_{0},I_{1})$ depending on the value of $c$. Clearly, $(J_{0},J_{1})$ depend on both $\theta^{B}_{\bar{I}_{t}}$ and $c$, but as Lemma 4.3 states, any correlation between $(J_{0},J_{1})$, $c$, and Alice’s information disappears if one does not have access to $\theta^{B}_{\bar{I}_{t}}$.
Lemma 4.3.
@@
A proof of both Lemmas 4.2 and 4.3 can be found in Appendix B.3. By setting $J=I_{t}$, Lemma 4.2 guarantees that Alice’s system’s state after the opening information has been sent is computationally indistinguishable from one that is completely uncorrelated with Bob’s measurement basis information in $\bar{I}_{t}$.
Additionally, by recalling that the value of $c$ is sampled independently of any of the considered systems, we know that the state $\rho_{{A^{\prime},\Theta^{A}_{\bar{I}_{t}},\Theta^{B}_{\bar{I}_{t}}},C}$ before $(J_{0},J_{1})$ is computed has the required product form and, from Lemma 4.3, we conclude that the state of all of Alice’s system at this point is computationally indistinguishable from a state uncorrelated with $C$. Let $\mathcal{E}$ be the operation Alice performs in her system from here to the end the protocol. By using Lemma A.2 (4) and grouping all of Alice’s systems into $S$, we obtain the desired result:
@@
4.3 Security against dishonest receiver
@@
We consider now the scenario in which Alice runs the protocol honestly and Bob runs an arbitrary program. For this analysis, note that Alice trusts her quantum state preparation and detection. We want to show that the state after finishing the protocol successfully satisfies Eq. (3). This means that the state at the end of the protocol can be described as a mixture of states where Bob’s system is uncorrelated with at least one of the two strings outputted by Alice. Similarly to the dishonest sender’s case, we assume that Bob’s laboratory consists of everything outside Alice’s, which means that he controls the communication channels and the environment. However, we do not assume that Bob is restricted to efficient computations.
@@
The values of Alice’s output strings depend on several quantities: Alice’s measurement outcomes, the choice of the $I_{t},J_{0},J_{1}$ subsets, and the choice of hashing function $f$ during the post-processing phase of the protocol. From all of these, the only ones that are not made explicitly public during the protocol’s execution are Alice’s measurement outcomes. Instead, partial information of these outcomes is revealed at different steps of the protocol. Let $x^{A}_{J_{0}},x^{A}_{J_{1}}$ be the sub-strings of measurement outcomes used to compute Alice’s outputs $m_{0},m_{1}$, respectively, and let $R$ denote Bob’s system at the end of the protocol (which includes all the systems that Alice sent during the execution of the protocol). In order to prove security we need to show that the joint state of the system $X^{A}_{J_{0}},X^{A}_{J_{1}},R$ can be written as a mixture of states $\rho^{b}$ (with $b\in\{0,1\}$) such that the conditional min-entropy $H_{\textnormal{min}}^{\varepsilon}(X^{A}_{J_{b}}|R)_{\rho^{b}}$ is high enough, so that we can use the leftover hash Lemma A.4 to guarantee that the outcome of the universal hashing $m_{b}=f(x^{A}_{J_{b}})$ is uncorrelated with $R$.
@@
At the start of the protocol the parties share a completely correlated entangled system. If the parties make measurements as intended, their outcomes will be only partially correlated, but if Bob was able to postpone his measurement until after Alice’s reveals her measurement bases, Bob could potentially obtain the whole information of $x^{A}$ by measuring in the appropriate basis on his system. To prevent this, Bob is required to commit his measurement bases and results to Alice before knowing which set is going to be tested. Then a statistical test is performed in Step (7) to estimate the correlation of Alice’s measurement outcomes with with the ones that Bob committed. As Lemma 4.4 states, any state passing the aforementioned test is such that, regardless of how Bob defines the sets $(J_{0},J_{1})$ during the string separation phase, there is a minimum of uncertainty that he has with respect to Alice’s measurement outcomes. Recall that, when Alice is honest, the overall state of the protocol before Step (8) will be a partially classical state, which could be written as a mixture over all of Alice’s classical information. Let $\vec{\tau}=(x^{A}_{I_{t}},\theta^{A},r,\textnormal{com},I_{t},I_{s},\textnormal{open}_{I_{t}})$ denote the transcript of the protocol up to Step (8), and let $\rho_{X^{A}B}(\vec{\tau},J_{0},J_{1})$ be the joint state of Alice’s measurement outcomes and Bob’s laboratory conditioned to $\vec{\tau},J_{0},J_{1}$.
Lemma 4.4.
@@
To reach the desired result, we will first show that a state $\tilde{\rho}_{T,\textnormal{SEP},X^{A},B}$ satisfying Lemma 4.4 (1) also satisfies a tighter version Lemma 3.2, and then use Lemma 4.4 (2) to attain the bound for the real protocol’s outcome. Since $\tilde{\rho}_{T,\textnormal{SEP},X^{A},B}$ is classical in both $T$ and SEP we can write the state of the joint system of Alice’s measurement outcomes and Bob’s laboratory as a mixture over all the possible transcripts at that point, that is:
@@
where $P(\vec{\tau},J_{0},J_{1})$ defines a probability distribution which is dependent on Bob’s behavior during the previous steps. We can now separate the $\tilde{\rho}_{X^{A}B}(\vec{\tau},J_{0},J_{1})$ in two categories depending on which of the $x^{A}_{J_{0}},x^{A}_{J_{1}}$ is the least correlated with Bob’s system. Consider the function $b(\vec{\tau},J_{0},J_{1})$ to be equal to 0 if $H_{\textnormal{min}}^{\varepsilon}(X^{A}_{J_{0}}|X^{A}_{J_{1}}B)_{\rho(\vec{\tau},J_{0},J_{1})}\geq H_{\textnormal{min}}^{\varepsilon}(X^{A}_{J_{1}}|X^{A}_{J_{0}}B)_{\rho(\vec{\tau},J_{0},J_{1})}$, and equal to $1$ otherwise. By regrouping the terms from (29) for which the value of $b$ is the same, we can rewrite the joint state as:
@@
where, from Lemma 4.4 and recalling that, as Lemma A.3 (5) states, the min-entropy of a mixture is lower bounded by that of the term with the least min-entropy, we know that
@@
where the second inequality follows from Lemma A.3 (3) and (4), and the max entropy term $H_{\max}(S_{b})$ is upper bounded by the size in bits of the syndrome $\ell=N_{\text{raw}}(f\cdot h(p_{\text{max}}+\delta_{1}))$. Now we can apply Equation (4.3) to Lemma A.4, which states that, for the outcomes $M_{0},M_{1}$ of the universal hashing by Alice in Step (12) and the Bob’s final system $R$ it holds that
@@
Finally, by applying Lemma A.1 (3) and (4) to Equations (30) and  (33), and then Lemma A.1 (1) to Eq. (28) we get the desired result:
@@
4.4 Composability considerations
Since OT protocols are mainly used as a subroutine of larger applications it is important to understand the composability properties of $\pi_{\textnormal{ROT}}$. In general, this is done through simulation-based composability frameworks. As mentioned in Section 1, this protocol is based on the BBSC construction, which has been proven secure in the quantum Universal Composability (UC) framework by Unruh [41] assuming access to an ideal commitment functionality. This means that we can understand the composability properties of $\pi_{\textnormal{ROT}}$ by understanding the respective properties of the underlying weakly-interactive commitment protocol.
It is well known that UC commitments are impossible to realize in the plain model [22, 25]. Because of this, protocols are often analyzed within a hybrid model, where the parties have access to some base external functionality. We show in Appendix 4.4 that there exists a family of commitment schemes that are both weakly-interactive and UC-secure in the classical access Random Oracle Model (ROM) [46]. This, in tandem with the aforementioned reduction of OT to commitments, results in the following theorem:
Theorem 4.1.
There exists a family of weakly-interactive commitment schemes in relation to which $\pi_{\textnormal{ROT}}$ is UC-secure in the classical access ROM.
In relation to Theorem 4.1, we want to emphasize that, even though limiting the access to the random oracle to be classical may seem at first strong in the context of a quantum protocol (where the parties are required access to some quantum capabilities), it has little impact in the resulting security of larger MPC protocols for which the security is analyzed in the classical setting.
Finally, we would like to stress the merits of Def. 2.1 by itself. In particular, this definition was studied in [47] and [44] and stated to ensure security when the protocol is executed sequentially. Furthermore, the indistinguishability properties stated in Def. 2.1 provide a very strong security guarantee and, because the protocol does not have external inputs and the indistinguishability relations include arbitrary external systems, these properties will still hold in any environment, which makes it relatively straightforward to analyze as part of bigger applications.
5 Experimental Implementation
5.1 Description of the Setup
@@
A schematic representation of the experimental setup can be seen in Fig. 6. Spontaneous parametric down conversion (SPDC), attributed to Alice, is used to create polarization entangled photon pairs in the state ${|{\Psi^{+}}\rangle}=\frac{1}{\sqrt{2}}\left({|{HH}\rangle}+{|{VV}\rangle}\right)$, which are coupled into optical fiber. One photon is sent through a 50/50 fiber beam splitter, probabilistically routing it to one of two polarization projection stages. There, a quarter-wave plate (QWP), a half-wave plate (HWP) and a polarizing beam splitter (PBS) are used to project the photons state onto the linear ($H/V$) or diagonal (+/+/-) basis, respectively. All photons are sent to superconducting nanowire single photon detectors (SNSPDs) and their arrival time is recorded using a time tagging module (TTM). The second photon of the state $|{\Psi^{+}}\rangle$}, attributed to Bob, travels through an equivalent probabilistic projection setup.
@@
We note that our entanglement-based implementation presents two main technological advantages over prepare-and-measure configurations:
  • It circumvents the need for a certified quantum random number generator or for classical pseudorandomness that may compromise the security of the quantum phase: instead of feeding random (or pseudorandom) sequences into the active polarization modulator of a prepare-and-measure scheme, the choice of BB84 state is performed in a passive and uniformly random way by the beamsplitters present in both Alice and Bob’s measurement setups (also known as "remote state preparation").
  • In free space, it avoids the need for active polarization modulation, which imposes a strict upper limit on the protocol’s repetition rate governed by the bandwidth of the Pockels Cell and its high-voltage amplifier, typically achieving a few hundred kHz to a few hundred MHz [49]. By generating entangled photons that are passively projected onto one of the four BB84 states instead, our OT rate is not limited by any active prepare-and-measure encoding routine, but only by our picosecond-pulsed pump rate of around $300$ MHz. With other SPDC sources reaching the GHz [50] to tens of GHz regimes [51], our passive state preparation routine can perform even better.
5.2 Practical protocol
The protocol is identical to that from $\pi_{\textnormal{ROT}}$ as described in Section 3, with the following amendments:
@@ 5.2 Practical protocol
      11 click: assign the correct measured bit value and report a successful round
@@
      22 clicks from the same basis: assign a random bit value to the measurement result and report a successful round
@@ 5.3 Practical security
Note that the statistical check performed by Alice in the second step of the amended protocol (Section 5.2) ensures security under the assumption that there is no coherence in the photon-number basis. This is the case in our implementation, since SPDC produces states of the form $\sum_{n=0}^{\infty}\sqrt{c_{n}}{|{n}\rangle}_{1}{|{n}\rangle}_{2}$ in the number basis $\{{|{n}\rangle}\}$ [53], leaving the individual subsystems in incoherent mixtures of the form $\sum_{n=0}^{\infty}c_{n}{|{n}\rangle}{\langle{n}|}$.
To account for the leakage caused by undetected multi-photon emissions to our OT rate expression, we effectively grant Bob an amount of information about Alice’s measurement outcomes equal to the number of indices in $I_{\bar{I}_{t}}$ associated to multi-photon events, upper bounded by $p_{\textnormal{multi}}(1-\alpha)N_{0}$ for large $N_{0}$. Subtracting this leak to the total entropy expression in Eq. (4.3) leads to a version of Lemma 3.2 for security against dishonest receiver corrected for the experimental implementation, which differs from the theoretical version by replacing Eq (13) with
@@
$\quad\quad+\frac{1}{2}\cdot 2^{\frac{1}{2}\left(n-N_{\textnormal{raw}}\left(\frac{1}{2}-\frac{2\delta_{2}}{1-2\delta_{2}}-h\left(\frac{p_{\textnormal{max}}+\delta_{1}}{\frac{1}{2}-\delta_{2}}\right)-f\cdot h(p_{\textnormal{max}}+\delta_{1})-\frac{p_{\textnormal{multi}}}{\frac{1}{2}-\delta_{2}}\right)\right)},$
6 Discussion